Equipment with an original cost of 75,000 and accumulated depreciation of 20,000 was sold at a loss of 7,000. As a result of this transaction, cash would
Answer #1

Book value of asset = Historical cost - Depreciation = 75000 - 20000 = 55000

Selling price = Book price - loss = 55000 - 7000 = 48000

Since asset is sold at 48000, cash received = $48000

Cash would increase by $48000